OpenStudy (toxicsugar22):

A block of metal has a volume of 14.0in3 and weighs 5.26lb . What is its density in grams per cubic centimeter?

OpenStudy (nurali):

1 pound = 453.59237 grams so 5.26 pounds = 2385.896 grams 1 inch cubed = 16.387064 centimeters cubed so 14 inches cubed= 229.418 cm cubed D=M/V D=2385.896/229.418 D=10.39 gram / cubic centimeter

OpenStudy (toxicsugar22):

10.39 is answer

OpenStudy (nurali):

Yes

OpenStudy (toxicsugar22):

get can yoou helpme with one more

OpenStudy (nurali):

i will try.

OpenStudy (toxicsugar22):

Acetone (fingernail-polish remover) has a density of 0.7857 g/cm3. What is the mass in grams of 17.52mL of acetone?

OpenStudy (nurali):

1 cm3 = 1 ml. Therefore, 17.52 mL has a mass of 0.7857g/mL x 17.52 mL which is 13.7g.

OpenStudy (nurali):

Yes

OpenStudy (nurali):

Yes

OpenStudy (toxicsugar22):

What is the volume in milliliters of 7.25g of acetone?

OpenStudy (toxicsugar22):

ok i will

OpenStudy (toxicsugar22):

but can you help me with that last question

OpenStudy (toxicsugar22):

What is the volume in milliliters of 7.25g of acetone?

OpenStudy (nurali):

V = mass/density = 7.25g / 0.7857 g/mL = 9.22 mL